Main INDEX
Monthly INDEX
PREV
NEXT
User name R. Feuerbach
Log entry time 06:27:29 on July24,2004
Entry number 129517
keyword=Q2 measurement
Run 1289 -- spot check and DAQ test. all okay
Asked for auto-gain on the BPMs, and after checks changed to spec DAQ, VDCs still off.
At 1µA, Hydrogen target in, the rates were recorded in run 1291 as:
T1=722kHz
T2=610kHz
T3=660kHz
T4=540kHz
Beam position for the past 12+hrs has been offset at BPMA/B (-0.16,-0.03) / (-0.33,-0.27). The first Q2 point is at this same position, the second is with a well-centered beam.
Reduced current to 0.2µA and turned off beam modulation, Rate check is run 1292. Turned on VDCs. Run 1294 is the spec-DAQ run.
Turned off VDCs and increased current to 1µA. Recorded beam position in run 1295, then had beam recentered to within .1mm of (0,0), beam position recorded in 1296. Turned on VDCs and took spec. DAQ run 1297.
Data in run 1294 and 1297 (VDCs, HAPPEX det, S2) look good.
Run
FIGURE 1